OpenStudy (anonymous):

how many atoms are there in .250 mole of C3H8

OpenStudy (australopithecus):

multiply by Avogadro number, it is a simple conversion 1mol = avagadros number thus (avagardos number (atoms)/1mol)(0.250mol) = # of atoms

OpenStudy (australopithecus):

notice mol cancels out and you are left with the unit atoms

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

\(N_A\frac{[\text{molecules}] }{[\text {mol}]}=\) avogadro's number of molecules \[ {N_A\frac{[\text{molecules}] }{[\text {mol}]}}\times0.250[\text {mol}][C_3H_8] =\# [\text{molecules}][C_3H_8] \] each molecule has Three carbon atoms and Eight Hydrogen atoms \[\frac{[\text{atoms}]}{ [\text{molecule}][C_3H_8] } =11\] \[\#[\text{molecules}][C_3H_8]* \frac{[\text{atoms}]}{ [\text{molecule}][C_3H_8] }= \# [ \text{atoms}] \]